OpenStudy (anonymous):

find constant C of x^2+11x+c (complete the square) in the xy plane has its vertex on the line y=x

OpenStudy (anonymous):

i gotta run but the idea is this vertex has first coordinate \[-\frac{11}{2}\] and second coordinate what you get when you plug that in since the vertex is on the line \[y=x\] you can set that result equal to \[-\frac{11}{2}\] and solve for c

OpenStudy (anonymous):

ty ^^

OpenStudy (mertsj):

c=99/4